A simple random sample of 40 items resulted in a sample mean of 60. The population standard deviation is σ =20 .

a. Compute the 95% confidence interval for the population mean. Round your answers to one decimal place.
( , )

b. Assume that the same sample mean was obtained from a sample of 130 items. Provide a 95% confidence interval for the population mean. Round your answers to two decimal places.
( , )
Answer #1

Solution:-

a) 95% confidence interval for the population mean is C.I = (53.80, 66.20).
C.I = \bar{x}\pm z_{\alpha /2}\ast \frac{\sigma }{\sqrt{n}}

C.I = 60 + 1.96*3.16228

C.I = 60 + 6.1981

C.I = (53.802 , 66.198)

b) 95% confidence interval for the population mean is C.I = (56.56, 63.44)

C.I = \bar{x}\pm z_{\alpha /2}\ast \frac{\sigma }{\sqrt{n}}

C.I = 60 + 1.96*1.75412

C.I = 60 + 3.4381

C.I = (56.562, 63.438)